# Quantum Reservoir Computing for Realized Volatility Forecasting

### v1

@@ IV.3 Feature selection for quantum reservoir computing
To see how increasing the features affects the prediction quality, in Fig. 7(a), we plot the realized volatility as well as the QR1 forecast for the first $1$, $4$, $7$, and $9$ features, which are shown in Fig. 6. As the figure clearly shows by increasing the features, the prediction improves until the number of features reach $n_{1}=7$. By further increasing the number of features, the prediction quality decreases as the number hidden qubits decreases and thus the reservoir lacks enough memory for incorporating the past information for future prediction. The corresponding results for the QR2 model is also shown in Fig. 7(b). The same conclusion about the impact of features on prediction quality can be deduced from the QR2 model. In addition, by comparing the corresponding panels in Fig. 7(a) Fig. 7(b) one can clearly see that the QR2 outperforms the QR1 prediction. This is expected as the the QR2 uses twice more measurement data for predicting the future outcomes.
IV.4 Model Interpretability
@@
To answer this we employ the Shapley value method, a model-agnostic technique grounded in cooperative game theory. Originally introduced in Ref. [119], the Shapley value offers a principled framework for fairly distributing the total payoff among participants in a coalition based on their individual contributions. This approach considers all possible combinations and permutations of features, helping to quantify the contribution of each feature to the model prediction. Lundberg and Lee [120] adapted this concept to machine learning, proposing Shapley values as a unified and interpretable measure of feature importance that satisfies key axioms such as consistency and local accuracy. In this spirit, our analysis applies Shapley values to the quantum reservoir computing framework to assess how macro-financial variables contribute to realized volatility forecasts, bridging interpretability with quantum-enhanced prediction. Here, we evaluate the Shapley values for the QR1 and QR2 models by a Julia package “ShapML” [121]. The choice of the features is decided according to one of following three strategies:
@@
    Feature-family contribution: In this strategy, we evaluate the contribution of each feature family irrespective of the time lag. Therefore, during the evaluation of the Shapley value when we evaluate the contribution of one feature type, e.g. for realized volatility $RV$, we consider contributions from all $\{RV_{t-3},RV_{t-2},RV_{t-1}\}$ time-lagged steps. The results are shown in Figs. 8(c) and (d) for the QR1 and QR2 models respectively. Interestingly, the Shapley value-based method furnishes a somewhat different ordering of feature importance compared to the forward selection method results depicted in Fig. 6. Note that this is not inconsistent and stems from the fact that the Shapley value is computed when all the features are embedded in the model while the forward selection method works by considering each individual feature in isolation.
@@
IV.5 Performance Metrics and Forecast Evaluation
@@
The final surviving set of models is denoted by ^$\widehat{\mathcal{M}}^{*}_{1-\alpha}$, the $(1-\alpha)$ MCS, which is asymptotically guaranteed to contain the model(s) with the lowest expected loss with probability at least $1-\alpha$. In this study, we set $\alpha=0.05$, ensuring that the resulting confidence set contains the best-performing model(s) with at least 95% probability in large samples. We complement the MCS analysis with pairwise Diebold-Mariano tests to assess the statistical significance of forecast performance differentials between models. This dual approach allows us to validate robustness in model rankings and mitigate the limitations associated with relying on a single evaluation metric.
@@
DM $\text{DM statistic}=\frac{\bar{d}}{\sqrt{\widehat{\sigma}^{2}/T}},$
@@
Table 3: Model Confidence Set. Performance of classical and quantum Models in predicting realized volatility (1997.08–2017.12). The MSE and QLIKE are presented alongside the MCS $p$-values ($P_{\text{MCS}}$). Asterisks () indicate inclusion in the model confidence set. The quantum reservoir computing models outperform all classical models, with QR2 exhibiting the best performance across all measures.
@@
MCS results.– The results of MCS are presented in Table 3. The table reports the performance metrics (MSE and QLIKE) along with the MCS $p$-values for classical time series models and quantum reservoir computing models. In particular, quantum models, especially QR2, demonstrate superior performance across all measures. From Table 3, we further observe that the QR2 model achieves the lowest MSE and QLIKE values, indicating its superior predictive accuracy. The MCS $p$-values further reinforce this finding, as QR2 attains a $p$-value of 1.0, signifying its inclusion in the superior set of models at the 95% significance level. In contrast, most classical models exhibit significantly higher loss values and lower MCS $p$-values, suggesting inferior performance. Examining the MCS $p$-values, models with asterisks are included in the Model Confidence Set $\mathcal{M}_{\text{MCS}}$, indicating that their predictive accuracy is statistically indistinguishable from the best-performing model. Among the classical models, only ARMAX, LSTMX, and RCX are included in the MCS based on their $p$-values, although their loss metrics are still higher than those of the quantum models. Additionally, QR1 is also included in the MCS, further highlighting the strong performance of quantum models. Examining the results, we also observe that including additional features sometimes enhances the model performance. For example, ARMAX and LSTMX exhibit lower loss values and higher MCS $p$-values compared to their counterparts AR3 and LSTM, suggesting that the inclusion of characteristics improves the accuracy of forecasting in these models. However, this improvement is not consistent across all models. Notably, HARX performs slightly worse than HAR in both MSE and QLIKE metrics, indicating that additional features do not always contribute to better performance. This suggests that the effectiveness of including features depends on the model structure and the relevance of the features.
@@ V Conclusions and Outlook
Acknowledgments. AB acknowledges support from the National Key R&D Program of China (Grant No. 2018YFA0306703), the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L).

### v2

@@ IV.4 Feature Selection for Quantum Reservoir Computing
To see how increasing the features affects the prediction quality, in Fig. 7(a), we plot the realized volatility as well as the QR1 forecast for the first 11, 44, 77, and 99 features, which are shown in Fig. 6. As the figure clearly shows by increasing the features, the prediction improves until the number of features reach $n_{1}=7$. By further increasing the number of features, the prediction quality cannot be further improved because the number of hidden qubits decreases; thus, the reservoir lacks sufficient memory to incorporate past information for future prediction. The corresponding results for the QR2 model is also shown in Fig. 7(b). The same conclusion about the impact of features on prediction quality can be deduced from the QR2 model. In addition, by comparing the corresponding panels in Fig. 7(a) Fig. 7(b) one can clearly see that the QR2 outperforms the QR1 prediction. This is expected as the the QR2 uses twice more measurement data for predicting the future outcomes.
IV.5 Model Interpretability
@@
To answer this we employ the Shapley value method, a model-agnostic technique grounded in cooperative game theory. Originally introduced in Ref. [120], the Shapley value offers a principled framework for fairly distributing the total payoff among participants in a coalition based on their individual contributions. This approach considers all possible combinations and permutations of features, helping to quantify the contribution of each feature to the model prediction. Lundberg and Lee [121] adapted this concept to machine learning, proposing Shapley values as a unified and interpretable measure of feature importance that satisfies key axioms such as consistency and local accuracy. In this spirit, our analysis applies Shapley values to the quantum reservoir computing framework to assess how macro-financial variables contribute to realized volatility forecasts, bridging interpretability with quantum-enhanced prediction. In theory, exact computation of Shapley values is exponential in the number of features; therefore, we typically rely on approximation methods in practice. In this work, we estimate Shapley values for the QR1 and QR2 models using the Monte Carlo sampling method [122], as implemented in the Julia package ShapML. The choice of the features is decided according to one of following three strategies:
@@
    Feature-family contribution: In this strategy, we evaluate the contribution of each feature family irrespective of the time lag. Therefore, during the evaluation of the Shapley value when we evaluate the contribution of one feature type, e.g. for realized volatility RVRV, we consider contributions from all $\{RV_{t-3},RV_{t-2},RV_{t-1}\}$ time-lagged steps. The results are shown in Figs. 8(c) and (d) for the QR1 and QR2 models respectively. Interestingly, the Shapley value-based method furnishes a somewhat different ordering of feature importance compared to the forward selection method results depicted in Fig. 6. Note that this is not inconsistent and stems from the fact that the Shapley value is computed when all the features are embedded in the model while the forward selection method works by considering each individual feature in isolation.
@@
IV.6 Performance Metrics and Forecast Evaluation
@@
The final surviving set of models is denoted by $\widehat{\mathcal{M}}^{*}_{1-\alpha}$, the $(1-\alpha)$ MCS, which is asymptotically guaranteed to contain the model(s) with the lowest expected loss with probability at least $1-\alpha$. In this study, we set $\alpha=0.05$, ensuring that the resulting confidence set contains the best-performing model(s) with at least 95% probability in large samples. We complement the MCS analysis with pairwise Diebold-Mariano tests to assess the statistical significance of forecast performance differentials between models. This dual approach allows us to validate robustness in model rankings and mitigate the limitations associated with relying on a single evaluation metric.
@@
$\text{DM statistic}=\frac{\bar{d}}{\sqrt{\widehat{\sigma}^{2}/T}},$
@@
Table 2: Model Confidence Set. Performance of classical and quantum Models in predicting realized volatility with predicting $S$-ahead step. The MSE and QLIKE are presented alongside the MCS $p$-values ($P_{\text{MCS}}$). Asterisks () indicate inclusion in the model confidence set. The quantum reservoir computing models outperform all classical models, with QR2 exhibiting the best performance across all measures.
S=1 (entire sample: 1997.08-2017.12) (1950.01- 2017.12)
@@
S=5 (entire sample: 1998.01-2017.12) (1950.01- 2017.12)
Model MSE QLike
Loss of MSE $P_{MCS}$ of MSE Loss of QLike $P_{MCS}$ of QLike
Classical Time Series Models
HAR 0.2143 $0.1291^{*}$ 2.9041 $0.1275^{*}$
HARX 0.2934 0.0044 4.5800 0.0040
AR1 0.2642 $0.0938^{*}$ 3.4136 $0.0937^{*}$
AR3 0.2134 $0.1302^{*}$ 2.8369 $0.1297^{*}$
ARMAX 0.2134 $0.1302^{*}$ 3.0703 $0.1297^{*}$
LSTM 0.1831 $0.1291^{*}$ 2.4600 $0.1275^{*}$
LSTMX 0.2200 $0.0925^{*}$ 3.4512 $0.0851^{*}$
RC 0.1528 $1.0000^{*}$ 2.0551 $1.0000^{*}$
RCX 0.1667 $0.6333^{*}$ 2.4605 $0.6251^{*}$
Quantum Reservoir Computing
QR1 0.1556 $0.7642^{*}$ 2.1518 $0.7703^{*}$
QR2 0.1663 $0.6333^{*}$ 2.2332 $0.6251^{*}$
@@
MCS results.– The results of MCS are presented in Table 2. The table reports the performance metrics (MSE and QLIKE) along with the MCS $p$-values for classical time series models and quantum reservoir computing models. In particular, when S=1, namely one-step ahead prediction, quantum models, especially QR2, demonstrate superior performance across all measures. From the upper table in Table 2, we further observe that the QR2 model achieves the lowest MSE and QLIKE values, indicating its superior predictive accuracy. The MCS $p$-values further reinforce this finding, as QR2 attains a $p$-value of 1.0, signifying its inclusion in the superior set of models at the 95% significance level. In contrast, most classical models exhibit significantly higher loss values and lower MCS $p$-values, suggesting inferior performance. Examining the MCS $p$-values, models with asterisks are included in the Model Confidence Set $\mathcal{M}_{\text{MCS}}$, indicating that their predictive accuracy is statistically indistinguishable from the best-performing model. Among the classical models, only ARMAX, LSTMX, and RCX are included in the MCS based on their $p$-values, although their loss metrics are still higher than those of the quantum models. Additionally, QR1 is also included in the MCS, further highlighting the strong performance of quantum models. Examining the results, we also observe that including additional features sometimes enhances the model performance. For example, ARMAX and LSTMX exhibit lower loss values and higher MCS $p$-values compared to their counterparts AR3 and LSTM, suggesting that the inclusion of characteristics improves the accuracy of forecasting in these models. However, this improvement is not consistent across all models. Notably, HARX performs slightly worse than HAR in both MSE and QLIKE metrics, indicating that additional features do not always contribute to better performance. This suggests that the effectiveness of including features depends on the model structure and the relevance of the features. When consider $S=5$, namely, the long-term prediction, both Quantum and Classical reservoir demonstrated outstanding performance compared with other models. The performances of QR1 and QR2 are very similar to that of the classic reservoir computing. Considering that we only used a simple reservoir model, these indirectly demonstrate the capabilities of Quantum reservoir.
@@ V Conclusions and Outlook
Quantum computing for machine learning still faces significant challenges, both in hardware and algorithmic development. On the hardware side, current quantum devices are still in their infancy, constrained by limited qubit counts, restricted connectivity, finite coherence times, and imperfect readout. Overcoming these limitations will require advances in error correction—a milestone likely still a decade away. Nevertheless, near-term quantum devices provide valuable opportunities to develop and test small-scale algorithms, paving the way for large-scale applications on future error-corrected quantum computers. From an algorithmic perspective, it remains unproven whether quantum computers can outperform classical machine learning methods on classical data. In this work, we demonstrate that quantum computers can indeed solve prediction tasks in stochastic time series. While our quantum reservoir learning approach shows advantages over several classical methods, we do not claim a definitive quantum advantage, as such a conclusion would require rigorous theoretical proof beyond the scope of this paper. Moreover, intrinsic limitations on QRC like the exponential concentration problem [129], whereby the predictions may become input-independent due to concentration of expectation values for many-qubit reservoirs, exist and have to be addressed in future work.
@@
Acknowledgments. AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528, No. 1251101297 and No. W2541020).
Appendix A Quantum Computation Preliminaries
Let us briefly provide an overview of the quantum theory necessary for understanding the subsequent sections of this paper. Quantum mechanics offers a mathematical framework originally developed to explain several experimental phenomena that challenged classical Newtonian physics in the late 19th and early 20th centuries. Readers interested in a more comprehensive exploration are encouraged to consult standard references such as Dirac [130] and Nielsen and Chuang [131]. Throughout this paper, we employ the Dirac notation commonly used in quantum physics. To assist readers from financial and econometric backgrounds, we present a minidictionary in Table 4, where the departures of the Dirac notation from the standard econometric notation as laid out in [132] are codified. In modern axiomatic form, quantum mechanics is completely described through the following four postulates.
Table 4: A mini-dictionary of Dirac notation for econometricians.
Object Econometric Notation Dirac Notation Comments
Scalar Variable $a$ $a$
Complex Conjugate of Scalar $\bar{a}$ or $^{*} aa^{*}$
Composite Vector $\textbf{{a}}\otimes\textbf{{b}}$ $|a\rangle\otimes|b\rangle\equiv|ab\rangle$ Produces a higher-dimensional vector
Complex Conjugate of Matrix A $\bar{\textbf{{A}}}$ $\mathbf{A}^{*}$ Element-wise complex conjugate
Adjugate of Matrix A $\textbf{{A}}^{\#}$ $\text{adj}(\mathbf{A})$
Determinant of Matrix A $\det(\textbf{{A}})$ $\det(\mathbf{A})$ Scalar value
Trace of Matrix A $\text{tr}(\textbf{{A}})$ $\operatorname{Tr}(\mathbf{A})$ Sum of diagonal elements
Identity Matrix $\textbf{{I}}_{n}$ $\mathbb{I}$
Zero Vector/Matrix 0 $\mathbf{0}$
Hermitian Operator $\textbf{{A}}=\textbf{{A}}^{\dagger}$ $A=A^{\dagger}$ Self-adjoint operator
Commutator of Operators AA and $B$ [A,B]=ABBA[A,B]=AB-BA [A,B]=ABBA[A,B]=AB-BA Measures non-commutativity
Anticommutator of Operators AA and $B$ $\{A,B\}=AB+BA$ $\{A,B\}=AB+BA$
Exponential of Operator $\exp(\textbf{{A}})$ $e^{A}$ Defined via power series
Fourier Transform of Function $f(x)$ $\mathcal{F}\{f(x)\}=\int_{-\infty}^{\infty}e^{-i\omega x}f(x)\,dx$ Same as econometric notation
Tensor Product of Matrices A and B $\textbf{{A}}\otimes\textbf{{B}}$ $A\otimes B$
Spin Operators N/A $S_{\alpha}=\frac{\hbar}{2}\alpha$ $\alpha=\{X,Y,Z\}$
Postulate 1: quantum states. – The complete information about a quantum system is encoded in its state vector $|\psi\rangle$, which is a ray of the system Hilbert space $\mathcal{H}$. The quantum bit (or qubit for short) is the smallest nontrivial unit in quantum computing, similar to the bit in classical computing. It is physically implemented by a two-level quantum system, such as the spin of an electron. Its mathematical form can be represented by a state vector in a Hilbert space $\mathcal{H}$, such that the single qubit state is defined as
And,
Postulate 2: composition of quantum systems. – When we consider a joint quantum system consisting of multiple qubits, the state of this system is described by the tensor product Hilbert space of each qubit $\mathcal{H}_{i}$. For example, if two qubits are described by their Hilbert spaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ respectively, the joint system comprising both these systems as subsystems is described by the tensor-product Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Thus, the state of two qubits is
$|\psi\rangle$ $=|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ (30)
[…]
The dimension of the $n$-qubit system is $2^{n}$, and hence the tensor product space is a $2^{n}$-dimensional complex vector space $\mathcal{C}^{2^{n}}$. The dimension of the $n$-qubit system increases exponentially in the number $n$ of the qubits. Quantum entanglement is another important property of quantum states that is different from the classical states. It occurs when a joint quantum system can not be represented as a tensor product of the subsystems that make up it. Mathematically, if two quantum systems AA and $B$ are entangled, the joint system ABAB might be represented as
$\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|,$ (34)
Postulate 3: quantum dynamics. – The time evolution of a closed quantum system describes how the state of the system evolves over time, which is the most important way to deal with the information. The closed means the evolution of systems that don’t interact with the rest of the world. The time evolution is governed by the Schrödinger equation, which is a fundamental partial differential equation in quantum mechanics.
The time-dependent Schrödinger equation is given by
$i\hbar\frac{d}{dt}|\psi(t)\rangle=H|\psi(t)\rangle,$ (35)
$|\psi(t)\rangle=e^{-\frac{i}{\hbar}Ht}|\psi(0)\rangle,$ (36)
where $|\psi(0)\rangle$ is the initial state of the system at time 0, and $e^{-\frac{i}{\hbar}Ht}$ is the time evolution operator.In practice, it is common to absorb the factor $\hbar$ into $H$, effectively setting $\hbar=1$. If the Hamiltonian $H$ depends on time, the situation is more complex, but we don’t use it in this paper. It is not difficult to see that the time evolution operator $e^{-\frac{i}{\hbar}Ht}$ is a unitary matrix, so this time evolution is a unitary transformation, which preserves the norm of the quantum states, meaning the probability conservation. For simplicity, we usually use $U(t)$ to represent $e^{-iHt}$ as
$|\psi(t)\rangle=U(t)|\psi(0)\rangle.$ (37)
$\rho(t)=U(t)\rho(0)U^{\dagger}(t),$ (38)
Postulate 4: quantum measurement. – Above, we described the evolution of a closed quantum system. However, to gain any information about the system, one has to necessarily consider an experimentalist as an external physical system observing it. Unlike classical mechanics, where in principle, such observations can be made without changing the observed system itself, quantum mechanics distinguishes itself by postulating that the quantum system undergoes irreversible change due to such a measurement. The general quantum measurement postulate is described by a collection measurement operator $\{M_{m}\}$. The index $m$ refers to the measurement outcomes that may occur in the experiment with the probability $p_{m}$ given by the Born rule $p_{m}=\langle\psi|M^{\dagger}_{m}M_{m}|\psi\rangle$, where $|\psi\rangle$ is the state being measured. To ensure the probabilities sum up to unity, the measurement operators satisfy the completeness equation
$\sum_{m}M_{m}^{\dagger}M_{m}=I$ (39)
In most cases of quantum computing, we care about the quantum measurement of an observable. An observable, i.e., a dynamical variable like position or momentum, is described by a self-adjoint operator AA (namely $A=A^{\dagger}$), acting on the system Hilbert space $\mathcal{H}$. Via the spectral theorem, any observable has a spectral decomposition
$|,A=\sum_{m}a_{m}\ket{a_{m}}\bra{a_{m}}$, (40)
[…]
$\langle A\rangle=\sum_{m}p_{m}a_{m}=\langle\psi|A|\psi\rangle$ (41)
Appendix B Numerical Simulation Method
[…]
In the partial-trace step, we obtain the reduced density matrix of the hidden state $\rho_{h}$ by tracing out subsystem II from the joint state $\rho_{Ih}$, i.e.,
$\rho_{h}=\mathrm{Tr}_{I}\!\left(\rho_{Ih}\right).$
Finally, the measurement outcome on each qubit is computed as the expectation value of the Pauli-$Z$ operator,
$\langle Z_{i}\rangle=\mathrm{Tr}\!\left(\rho_{Ih}Z_{i}\right).$
Appendix C Exponential concentration of Quantum Measurement
In this article[110], the authors discuss exponential concentration: as the circuit depth increases, the system size increases, entanglement increases, or the noise strength increases, the measurement expectation values converge exponentially to a variable-independent value, such that the number of measurement shots growths exponentially to identify different expectation values. In our paper, however, we consider a specific example—namely, the application of QRC to predicting RV. We use a fixed system size (the number of qubits is 10) and circuit depth (3 layers). By calculating the variance of the expectation value on each qubit, we find that it does not exhibit exponential concentration (see Fig. 9), thereby demonstrating the feasibility of QRC for this problem.
Figure 9: Variance of the measured expectation values across different qubits (left: QR1 model; right: QR2 model).
Appendix D The hyperparameters of Classical Reservoir computing and LSTM
For the LSTM(x), we used stack LSTM model with 2 layers. The optimizer is ADAM, with 0.001 learning rate. The number of iteration is 100 epochs. And the batch size is 64. After these hyperparameters are chosen. We train these LSTM with different hidden size from 10 to 200 with 10 interval. We get for the LSTM model, the hidden size is 60 achieving best result, and LSTMX model, 50 achieving bset performance (see Fig. 10 right).
Figure 10: The MSE of RCx (left) and LSTMx (right) as the hidden state size increases.
For Classical reservoir computing, RC(X), Apart from the number of neurons in reservoir, the other key hyperparameters are the leak rate (lr), controlling the time constant of the reservoir, the spectral radius of W (sr), the maximum absolute eigenvalue of the reservoir, and the input scaling (is), the coefficient applied on $W_{in}$. We set them as $lr=0.6$, $sr=0.9$, $is=0.1$ by trying a lot of configurations. This document provides a detailed description of these hyperparameters. Based on these hyperparameters, We train RC(X) models with the number of neurons varying from 10 to 200 with 10 interval. We get for the RC model, the number of neurons is 50 achieving best result, and RCx model, 20 achieving the best performance (see Fig. 10 left). As shown in Fig. 10 depicted, increasing the number of hidden nodes (neurons) does not always reduce errors. So we selected the best model by comparing the performance of different hidden nodes (neurons).